A monopolist with the cost function C=1/2Q2 faces a demand curve Q=12-P,

A.  What will be his equilibrium price and quantity?

B.  If for some reason the firm behaves as if it were in a perfectly competitive industry, what will equilibrium price and quantity?

C.  How much money will the firm require to forgo monopoly profits and behave competitively instead? 



1
Expert's answer
2022-02-16T08:28:11-0500

A. His equilibrium price and quantity are:

MR = MC,

MR = TR'(Q) = 12 - 2Q,

MC = TC'(Q) = Q,

12 - 2Q = Q,

Q = 4 units,

P = 12 - 4 = 8.

B. If for some reason the firm behaves as if it were in a perfectly competitive industry, then the equilibrium price will be lower and quantity higher.

C. The firm will require to forgo some part of monopoly profits to behave competitively instead at the point MC = D.